# Crosscap states and duality of Ising field theory in two dimensions

### v1

@@ Abstract
(±$(\pm i)^{n}c_{j_{1}}^{\dagger}c_{j_{1}+N/2}^{\dagger}\cdots c_{j_{n}}^{\dagger}c_{j_{n}+N/2}^{\dagger}=\begin{cases}c_{j_{1}}^{\dagger}\cdots c_{j_{n}}^{\dagger}c_{j_{1}+N/2}^{\dagger}\cdots c_{j_{n}+N/2}^{\dagger}&\quad n\;\mathrm{even}\\ \pm i\;c_{j_{1}}^{\dagger}\cdots c_{j_{n}}^{\dagger}c_{j_{1}+N/2}^{\dagger}\cdots c_{j_{n}+N/2}^{\dagger}&\quad n\;\mathrm{odd}\\ \end{cases}\,,$ (S37)
@@
C.3 C. Lattice crosscap overlaps for excited states
@@
=$=\begin{cases}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\cdot\langle\psi_{0}(h)|B_{\mathrm{latt}}^{\pm}\rangle&\quad M\;\mathrm{even}\\ \mp(-1)^{N/2}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\cdot\langle\psi_{0}(h)|B_{\mathrm{latt}}^{\pm}\rangle&\quad M\;\mathrm{odd}\\ \end{cases}\,.$ (S45)
@@
$=i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\cdot(-1)^{N/2}\left[\cos\left(\frac{\pi}{4}-(-1)^{N/2}\Theta\right)-\sin\left(\frac{\pi}{4}-(-1)^{N/2}\Theta\right)\right]$
$=i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\cdot(-1)^{N/2}\sqrt{2}\sin\left[(-1)^{N/2}\Theta\right]$
@@
The crosscap overlaps of the other lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle$ [Eq. (S32)] can be obtained via the Kramers-Wannier duality. Using Eqs. (S18) and (S27), we find that the eigenstate $|\psi_{k_{1}\cdots k_{M}}(h)\rangle$ transforms under the Kramers-Wannier duality transformation as
@@
In summary, the lattice crosscap overlaps for $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$ are given by
$\langle\psi_{k_{1}\cdots k_{M}}(h)|\mathcal{C}_{\mathrm{latt}}^{+}\rangle=(-1)^{M}\langle\psi_{k_{1}\cdots k_{M}}(1/h)|\mathcal{C}_{\mathrm{latt}}^{-}\rangle=\begin{cases}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{2}\cos\Theta(h)&\quad M\;\mathrm{even}\\ i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{2}\sin\Theta(h)&\quad M\;\mathrm{odd}\\ \end{cases}\,.$ (S50)
Specifically, at the critical point $h=1$, we have $\theta_{k}=\frac{k}{2}$ [Eq. (S6)] and the angle variable $\Theta(h=1)$ [Eq. (S42)] is
$\Theta(h=1)=\frac{\pi}{4}+\frac{1}{4}\sum_{k>0}(-1)^{n_{k}}k=\frac{\pi}{8}\,.$ (S51)
from which we obtain the non-vanishing lattice crosscap overlaps
$\langle\psi_{k_{1}\cdots k_{M}}(h=1)|\mathcal{C}_{\mathrm{latt}}^{+}\rangle=(-1)^{M}\langle\psi_{k_{1}\cdots k_{M}}(h=1)|\mathcal{C}_{\mathrm{latt}}^{-}\rangle=\begin{cases}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{\frac{2+\sqrt{2}}{2}}&\quad M\;\mathrm{even}\\ i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{\frac{2-\sqrt{2}}{2}}&\quad M\;\mathrm{odd}\\ \end{cases}\,,$ (S52)
which corresponds to Eq.(8) in the main text.
Therefore, the lattice crosscap states $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$ are expanded in the eigenbasis of the critical Ising chain as
@@
$+\frac{e^{-i\pi/8}}{\sqrt{2}}\exp\left[\mp i\sum_{k>0}(-1)^{n_{k}}d_{-k}^{\dagger}(h=1)d_{k}^{\dagger}(h=1)\right]|\psi_{0}(h=1)\rangle\,.$ (S53)
Appendix D IV. Ising CFT and conformal crosscap states
@@
The critical Ising chain in the continuum limit is described by the Ising CFT. In this section, we identify the continuum counterparts of the lattice crosscap states $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$, denoted as $|\mathcal{C}_{\pm}\rangle$, and verify that $|\mathcal{C}_{\pm}\rangle$ are conformal crosscap states in the Ising CFT.
D.1 A. Ising CFT and its Majorana free field representation
We first briefly review the operator formalism of the Ising CFT, where we restrict ourselves in the NS sector for our purpose.
@@
The Ising CFT can be formulated with the free Majorana Hamiltonian:
$H^{\mathrm{NS}}_{0}=\frac{i}{4\pi}\int_{0}^{L}\mathrm{d}x\left[\chi\partial_{x}\chi-\bar{\chi}\partial_{x}\bar{\chi}\right]\,.$ (S54)
@@
$\chi$ and $\bar{\chi}$ are Majorana fields, with the mode expansion:
@@
We introduce the normal-ordered stress tensors
$:T:(z)=-\frac{1}{2}:\chi(z)\partial\chi(z):\,,\quad:\bar{T}:(\bar{z})=-\frac{1}{2}:\bar{\chi}(\bar{z})\bar{\partial}\bar{\chi}(\bar{z}):\,,$ (S56)
where $\partial=\frac{1}{2}(\partial_{\tau}-i\partial_{x})$ and ¯$\bar{\partial}=\frac{1}{2}(\partial_{\tau}+i\partial_{x})$ are complex derivatives, and ::$\cdots$:: ⋯ : denotes normal ordering. The mode expansions of the stress tensors are given by
$:T:(z)=\left(\frac{2\pi}{L}\right)^{2}\sum_{n\in\mathbb{Z}}L_{n}e^{-\frac{2\pi}{L}nz}\,,\quad:\bar{T}:(\bar{z})=\left(\frac{2\pi}{L}\right)^{2}\sum_{n\in\mathbb{Z}}\bar{L}_{n}e^{-\frac{2\pi}{L}n\bar{z}}\,,$ (S57)
where
$L_{n}=\frac{1}{2}\sum_{m\in\mathbb{Z}}(m-1/2):b_{n-m+1/2}b_{m-1/2}:\,,\quad\bar{L}_{n}=\frac{1}{2}\sum_{m\in\mathbb{Z}}(m-1/2):\bar{b}_{n-m+1/2}\bar{b}_{m-1/2}:$ (S58)
are the Virasoro generators of the Ising CFT.
In terms of the Virasoro generators, the Hamiltonian becomes
$H^{\mathrm{NS}}_{0}$ $=\frac{1}{2\pi}\int_{0}^{L}\mathrm{d}x\left[:T:(ix)+:\bar{T}:(-ix)-\left(\frac{2\pi}{L}\right)^{2}\frac{c}{12}\right]$
$=\frac{2\pi}{L}\left[L_{0}+\bar{L}_{0}-\frac{c}{12}\right]$
$=\frac{2\pi}{L}\left[\sum_{n=1}^{\infty}(n-\frac{1}{2})\left(b_{-n+\frac{1}{2}}b_{n-\frac{1}{2}}+\bar{b}_{-n+\frac{1}{2}}\bar{b}_{n-\frac{1}{2}}\right)-\frac{c}{12}\right]\,,~{}$ (S59)
where $c=1/2$ is the central charge.
The primary state $|\varepsilon,\bar{\varepsilon}\rangle$, corresponding to the energy field $\varepsilon$, is obtained via the state-operator correspondence:
$|\varepsilon,\bar{\varepsilon}\rangle=\lim_{|w|\to 0}\varepsilon(w,\bar{w})|0\rangle_{\mathrm{NS}}=-ib_{-1/2}\bar{b}_{-1/2}|0\rangle_{\mathrm{NS}}\,,$ (S60)
where $|0\rangle_{\mathrm{NS}}\equiv|1,\bar{1}\rangle$, denoting the ground state of the Ising CFT, is the primary state corresponding to the identity field $1$. The fermionic representation of the energy field reads $\varepsilon(w,\bar{w})=-i\chi(w)\bar{\chi}(\bar{w})$, where $\chi(w)=(\frac{2\pi}{L}w)^{-1/2}\chi(z)$ and $\bar{\chi}(\bar{w})=(\frac{2\pi}{L}\bar{w})^{-1/2}\bar{\chi}(\bar{z})$ are the Majorana fields on the plane, with $w=e^{\frac{2\pi}{L}z}$.
Any eigenstate in the NS sector is generated by acting the Varasoro generators on the top of the primary states $|1,\bar{1}\rangle$ or $|\varepsilon,\bar{\varepsilon}\rangle$:
$\prod_{n=1}^{\infty}L_{-n}^{p_{n}}\bar{L}_{-n}^{\bar{p}_{n}}|\alpha,\bar{\alpha}\rangle\,,\quad\alpha=1\;\text{or}\;\varepsilon\,,$ (S61)
where $p_{n}$ and $\bar{p}_{n}$ are non-negative integers. However, some states (as well as certain linear combinations) in Eq. (S61) are zero, known as null states.
D.2 B. Ising conformal crosscap states
The Ising CFT [Eq. (S59)] is the effective theory of the critical Ising chain [Eq. (S1) with $h=1$], where the Bogouliubov modes near ±$\pm\pi$ are identified with the momentum modes of the Majorana fields:
@@
$|B_{\pm}\rangle=\exp\left[\pm\sum_{n=1}^{\infty}(-1)^{n}b_{-n+\frac{1}{2}}\bar{b}_{-n+\frac{1}{2}}\right]|0\rangle_{\mathrm{NS}}\,.~{}$ (S65)
@@
A state $|\mathcal{C}\rangle$ is termed as the conformal crosscap state if it satisfies the sewing condition:
$\left[L_{n}-(-1)^{n}\bar{L}_{-n}\right]|\mathcal{C}\rangle=0\,,\quad\forall n\in\mathbb{Z}\,.$ (S66)
Using the fermionic representation of the Virasoro generators $L_{n}$ and $\bar{L}_{n}$ given by Eq. (S58), we can directly verify that $|\mathcal{C}_{\pm}\rangle$ [Eq. (S64)] satisfy the sewing condition in Eq. (S66). Thus, $|\mathcal{C}_{\pm}\rangle$ are conformal crosscap states of the Ising CFT.
Let us verify this point more explicitly by considering the following crosscap Ishibashi states [21]:
$|\alpha\rangle\rangle_{\mathcal{C}}=\prod_{n=1}^{\infty}\left[\sum_{p_{n}=0}^{\infty}(-1)^{np_{n}}L_{-n}^{p_{n}}\bar{L}_{-n}^{p_{n}}\right]|\alpha,\bar{\alpha}\rangle=(-1)^{L_{0}-h_{\alpha}}\prod_{n=1}^{\infty}\left[\sum_{p_{n}=0}^{\infty}L_{-n}^{p_{n}}\bar{L}_{-n}^{p_{n}}\right]|\alpha,\bar{\alpha}\rangle\,,\quad\alpha=1,\varepsilon\,,$ (S67)
where $h_{\alpha}$ is the conformal weight of the primary field. Labeled by the primary field $\alpha=1,\varepsilon$, the crosscap Ishibashi state $|\alpha\rangle\rangle_{\mathcal{C}}$ is the conformal crosscap state supported on the Verma module of the primary state $|\alpha,\bar{\alpha}\rangle$.
Utilizing the fermionic representation of the Virasoro generators [Eq. (S58)], we can expand the crosscap Ishibashi states in the fermionic basis:
$|1\rangle\rangle_{\mathcal{C}}$ $=\sum_{M\;\mathrm{even}}\sum_{0<n_{1}<\cdots<n_{M}}(-1)^{\sum_{\alpha=1}^{M}(n_{\alpha}-1/2)}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\prod_{\alpha=1}^{M}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}\,,$
$|\varepsilon\rangle\rangle_{\mathcal{C}}$ $=-i\sum_{M\;\mathrm{odd}}\sum_{0<n_{1}<\cdots<n_{M}}(-1)^{-h_{\varepsilon}+\sum_{\alpha=1}^{M}(n_{\alpha}-1/2)}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\prod_{\alpha=1}^{M}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}\,,$ (S68)
where we used the fermionic representation of the primary state $|\varepsilon,\bar{\varepsilon}\rangle=-ib_{-1/2}\bar{b}_{-1/2}|0\rangle_{\mathrm{NS}}$ [Eq. (S60)].
The crosscap Ishibashi states $|1\rangle\rangle_{\mathcal{C}}$ and $|\varepsilon\rangle\rangle_{\mathcal{C}}$ can be expressed in terms of $|B_{\pm}\rangle$ [Eq. (S65)]:
$|1\rangle\rangle_{\mathcal{C}}$ $=\sum_{M\;\mathrm{even}}\sum_{0<n_{1}\cdots<n_{M}}(-1)^{\sum_{\alpha=1}^{M}(n_{\alpha}-1/2)}\cdot(-1)^{M/2}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}$
@@
$|\varepsilon\rangle\rangle_{\mathcal{C}}$ $=-i\sum_{M\;\mathrm{odd}}\sum_{0<n_{1}<\cdots<n_{M}}(-1)^{-1/2+\sum_{\alpha=1}^{M}(n_{\alpha}-1/2)}\cdot(-1)^{\frac{M-1}{2}}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}$

### v2

@@ Abstract
$(\pm i)^{n}c_{j_{1}}^{\dagger}c_{j_{1}+N/2}^{\dagger}\cdots c_{j_{n}}^{\dagger}c_{j_{n}+N/2}^{\dagger}=\begin{cases}c_{j_{1}}^{\dagger}\cdots c_{j_{n}}^{\dagger}c_{j_{1}+N/2}^{\dagger}\cdots c_{j_{n}+N/2}^{\dagger}&\quad n\;\mathrm{even}\\ \pm i\;c_{j_{1}}^{\dagger}\cdots c_{j_{n}}^{\dagger}c_{j_{1}+N/2}^{\dagger}\cdots c_{j_{n}+N/2}^{\dagger}&\quad n\;\mathrm{odd}\end{cases},$ (S47)
@@
C.3 C. Lattice crosscap overlaps for the excited states
@@
$=\begin{cases}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\cdot\langle\psi_{0}(h)|B_{\mathrm{latt}}^{\pm}\rangle&\quad M\;\mathrm{even}\\ \mp(-1)^{N/2}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\cdot\langle\psi_{0}(h)|B_{\mathrm{latt}}^{\pm}\rangle&\quad M\;\mathrm{odd}\end{cases}.$ (S55)
@@
$=i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\,(-1)^{N/2}\sqrt{2}\sin\left[(-1)^{N/2}\Theta\right]$
@@
In summary, the lattice crosscap overlaps for $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$ are given by
$\langle\psi_{k_{1}\cdots k_{M}}(h)|\mathcal{C}_{\mathrm{latt}}^{+}\rangle=\begin{cases}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{2}\cos\Theta(h)&\quad M\;\mathrm{even}\\ i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{2}\sin\Theta(h)&\quad M\;\mathrm{odd}\end{cases}.$ (S58)
C.4 D. Overlaps for the dual and modified lattice crosscap states
The crosscap overlaps of the dual lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle$ [Eq. (S32)] can be obtained via the Kramers-Wannier duality. Using Eqs. (S18) and (S27), we find that the eigenstate $|\psi_{k_{1}\cdots k_{M}}(h)\rangle$ transforms under the Kramers-Wannier duality transformation as
@@
For the modified lattice crosscap state $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$, we have the analogous fermionic representation:
$(-1)^{N/2}|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle=\frac{1+i}{2}|B_{\mathrm{latt}}^{+}\rangle+\frac{1-i}{2}|B_{\mathrm{latt}}^{-}\rangle\,$, (S61)
where $|B_{\mathrm{latt}}^{\pm}\rangle$ are the fermionic Gaussian states defined in Eq. (S46). This enables analytic computation of overlaps of $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$ with TFIC eigenstates:
$\langle\psi_{k_{1}\cdots k_{M}}(h)|\mathcal{C}_{\mathrm{latt}}^{\prime}\rangle=\begin{cases}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{2}\sin\Theta(h)&\quad M\;\mathrm{even}\\ -i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{2}\cos\Theta(h)&\quad M\;\mathrm{odd}\end{cases}.$ (S62)
C.5 E. Expansion of crosscap states in the critical Ising chain eigenbasis
Specifically, at the critical point of TFIC [Eq. (S1) with $h=1$], we have $\theta_{k}=\frac{k}{2}$ [Eq. (S6)] and the angle variable $\Theta(h=1)$ [Eq. (S52)] is
$\Theta(h=1)=\frac{\pi}{4}+\frac{1}{4}\sum_{k>0}(-1)^{n_{k}}k=\frac{\pi}{8}\,$, (S63)
from which we obtain the non-vanishing overlaps for the lattice crosscap states
$\langle\psi_{k_{1}\cdots k_{M}}(h=1)|\mathcal{C}_{\mathrm{latt}}^{+}\rangle=(-1)^{M}\langle\psi_{k_{1}\cdots k_{M}}(h=1)|\mathcal{C}_{\mathrm{latt}}^{-}\rangle=\begin{cases}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{\frac{2+\sqrt{2}}{2}}&\quad M\;\mathrm{even}\\ i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{\frac{2-\sqrt{2}}{2}}&\quad M\;\mathrm{odd}\end{cases},$ (S64)
and for the modified crosscap state
$\langle\psi_{k_{1}\cdots k_{M}}(h=1)|\mathcal{C}_{\mathrm{latt}}^{\prime}\rangle=\begin{cases}(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{\frac{2-\sqrt{2}}{2}}&\quad M\;\mathrm{even}\\ -i(-1)^{\sum_{\alpha=1}^{M}n_{k_{\alpha}}}\sqrt{\frac{2+\sqrt{2}}{2}}&\quad M\;\mathrm{odd}\end{cases}.$ (S65)
Note that Eq. (S64) corresponds to Eq. (10) in the main text.
According to Eqs. (S64) and  (S65), the lattice crosscap states $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$ are expanded in the eigenbasis of the critical Ising chain as
@@
$\phantom{=}\;+\frac{e^{-i\pi/8}}{\sqrt{2}}\exp\left[\mp i\sum_{k>0}(-1)^{n_{k}}d_{-k}^{\dagger}(h=1)d_{k}^{\dagger}(h=1)\right]|\psi_{0}(h=1)\rangle\,$, (S66)
and the modified crosscap state $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$ is expanded as
@@
$|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$ $=\frac{-ie^{i\pi/8}}{\sqrt{2}}\exp\left[i\sum_{k>0}(-1)^{n_{k}}d_{-k}^{\dagger}(h=1)d_{k}^{\dagger}(h=1)\right]|\psi_{0}(h=1)\rangle$
$\phantom{=}\;+\frac{ie^{-i\pi/8}}{\sqrt{2}}\exp\left[-i\sum_{k>0}(-1)^{n_{k}}d_{-k}^{\dagger}(h=1)d_{k}^{\dagger}(h=1)\right]|\psi_{0}(h=1)\rangle\,.$ (S67)
Appendix D IV. Ising conformal crosscap states
@@
In the continuum limit, the critical Ising chain is described by the Ising CFT. In this section, we first review the Pradisi-Sagnotti-Stanev (PSS) conformal crosscap states in the Ising CFT and then derive the associated dual state via the Kramers-Wannier topological defect line (TDL). Using the Majorana free field representation, we identify the continuum counterparts of the lattice crosscap states $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$, denoted as $|\mathcal{C}_{\pm}\rangle$, and determine their relation to the PSS crosscap states.
D.1 A. Pradisi-Sagnotti-Stanev crosscap states
The Ising CFT is the simplest example of the rational CFT with charge conjugation modular invariant. In such theories, conformal crosscap states can be systematically constructed from the bulk conformal data.
Let the modular $\mathcal{S}$-matrix be $S_{ab}$ and the $\mathcal{T}$-matrix be $T_{ab}=\delta_{ab}\,e^{2\pi i(h_{a}-\tfrac{c}{24})}$, where $h_{a}$ is the conformal weight of the chiral primary field, and $c$ is the central charge. The so-called PSS crosscap states can then be formally defined, each labeled by a primary field $a$ [22, 23, 35],
$|\mathcal{C}_{a}\rangle=\sum_{b}\frac{P_{ab}}{\sqrt{S_{0b}}}\,|b\rangle\rangle_{\mathcal{C}}\,$, (S68)
where the $P_{ab}$ is a symmetric and unitary matrix, defined as
$P=\sqrt{T}\,S\,T^{2}\,S\,\sqrt{T}\,$, (S69)
and $|a\rangle\rangle_{\mathcal{C}}\equiv e^{i\pi(L_{0}-h_{a})}|a\rangle\rangle$ are crosscap Ishibashi states, with $L_{0}$ the zeroth Virasoro generator and $|a\rangle\rangle$ the usual boundary CFT Ishibashi states. If the primary field $a$ is a simple current of the theory, the corresponding PSS crosscap state $|\mathcal{C}_{a}\rangle$ is physical [35]. In particular, the identity field is a trivial simple current, and the associated state $|\mathcal{C}_{0}\rangle$ is known as the “standard” PSS crosscap state.
Verlinde TDLs are defined as [65]
$\mathcal{D}_{a}=\sum_{b}\frac{S_{ab}}{S_{0b}}\mathcal{P}_{b}\,$, (S70)
where $\mathcal{P}_{b}$ projects onto the sector labeled by the primary field $b$. Acting with $\mathcal{D}_{a}$ on a PSS crosscap state yields
$\mathcal{D}_{a}|\mathcal{C}_{b}\rangle=\sum_{c}Y_{ab}^{c}\,|\mathcal{C}_{c}\rangle\,$, (S71)
where
$Y_{ab}^{c}\equiv(Y_{a})_{bc}=\sum_{d}\frac{S_{ad}\,P_{bd}\,P^{*}_{cd}}{S_{0d}}$ (S72)
is an integer-valued tensor [23, 66, 35, 67].
The Ising CFT, with central charge $c=\frac{1}{2}$, has three (chiral) primary fields $\mathbbm{1}\,,\varepsilon\,,\sigma$, with conformal weights $0\,,\frac{1}{2}\,,\frac{1}{16}$, respectively. Its modular $\mathcal{S}$-matrix is
$S=\frac{1}{2}\begin{pmatrix}1&1&\sqrt{2}\\ 1&1&-\sqrt{2}\\ \sqrt{2}&-\sqrt{2}&0\end{pmatrix}\,$, (S73)
from which the $P$-matrix and the $Y$-tensor can be computed explicitly:
$P=\begin{pmatrix}\cos(\pi/8)&\sin(\pi/8)&0\\ \sin(\pi/8)&-\cos(\pi/8)&0\\ 0&0&1\end{pmatrix}\,,\quad Y_{\mathbbm{1}}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\,,\quad Y_{\varepsilon}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix}\,,\quad Y_{\sigma}=\begin{pmatrix}1&1&0\\ 1&-1&0\\ 0&0&0\end{pmatrix}\,$. (S74)
From Eq. (S68), two PSS crosscap states, corresponding to the simple currents $\mathbbm{1}$ and $\varepsilon$, are given by [22, 23]
$|\mathcal{C}_{\mathbbm{1}}\rangle$ $=\sqrt{\frac{2+\sqrt{2}}{2}}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}+\sqrt{\frac{2-\sqrt{2}}{2}}|\varepsilon\rangle\rangle_{\mathcal{C}}\,$,
$|\mathcal{C}_{\varepsilon}\rangle$ $=\sqrt{\frac{2-\sqrt{2}}{2}}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}-\sqrt{\frac{2+\sqrt{2}}{2}}|\varepsilon\rangle\rangle_{\mathcal{C}}\,.$ (S75)
Although these two PSS crosscap states are both physical, their crosscap partition functions $\langle\mathcal{C}_{a}|e^{-\beta H_{0}}|\mathcal{C}_{a}\rangle$ ($a=\mathbbm{1},\varepsilon$) differ. Under a modular-$\mathcal{S}$ transformation, the one for $|\mathcal{C}_{\mathbbm{1}}\rangle$ gives the standard Klein bottle partition function, while that for $|\mathcal{C}_{\varepsilon}\rangle$ yields the non-standard Klein bottle partition function [35].
The Verlinde line $\mathcal{D}_{\sigma}=\sqrt{2}(\mathcal{P}_{\mathbbm{1}}-\mathcal{P}_{\varepsilon})$ realizes Kramers-Wannier duality in the Ising CFT. Acting with $\mathcal{D}_{\sigma}$ on the (standard) PSS crosscap state $|\mathcal{C}_{\mathbbm{1}}\rangle$ yields a dual conformal crosscap state, which by Eq. (S71) is a superposition of $|\mathcal{C}_{\mathbbm{1}}\rangle$ and $|\mathcal{C}_{\varepsilon}\rangle$ [36]:
$\frac{\mathcal{D}_{\sigma}}{\langle\mathcal{D}_{\sigma}\rangle}|\mathcal{C}_{\mathbbm{1}}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}-\sqrt{\frac{2-\sqrt{2}}{2}}|\varepsilon\rangle\rangle_{\mathcal{C}}=\frac{1}{\sqrt{2}}\left(|\mathcal{C}_{\mathbbm{1}}\rangle+|\mathcal{C}_{\varepsilon}\rangle\right)\,$, (S76)
where $\langle\mathcal{D}_{\sigma}\rangle=\sqrt{2}$ denotes the vacuum expectation value of the TDL. The crosscap states $|\mathcal{C}_{\mathbbm{1}}\rangle$ and $\frac{\mathcal{D}_{\sigma}}{\langle\mathcal{D}_{\sigma}\rangle}|\mathcal{C}_{\mathbbm{1}}\rangle$ share the same crosscap partition function and are mapped into each other under the Kramers-Wannier duality transformation.
D.2 B. Majorana free field representation
@@
The operator formalism of the Ising CFT can be formulated in terms of the free Majorana Hamiltonian:
$H^{\mathrm{NS}}_{0}=\frac{i}{4\pi}\int_{0}^{L}\mathrm{d}x\left[\chi\partial_{x}\chi-\bar{\chi}\partial_{x}\bar{\chi}\right]\,$, (S77)
@@
where we restrict ourselves in the NS sector for our purpose. Here $\chi$ and $\bar{\chi}$ are Majorana fields, with the mode expansion:
@@
The Virasoro generators can be expressed in terms of these modes as
$L_{n}=\frac{1}{2}\sum_{m\in\mathbb{Z}}(m-1/2):b_{n-m+1/2}b_{m-1/2}:\,,\quad\bar{L}_{n}=\frac{1}{2}\sum_{m\in\mathbb{Z}}(m-1/2):\bar{b}_{n-m+1/2}\bar{b}_{m-1/2}:\,$, (S79)
where $\cdots:$ denotes normal ordering. The primary states $|\mathbbm{1}\rangle$ and $|\varepsilon\rangle$ are
$|\mathbbm{1}\rangle=|0\rangle_{\mathrm{NS}}\,,\quad|\varepsilon\rangle=-ib_{-1/2}\bar{b}_{-1/2}|0\rangle_{\mathrm{NS}}\,$, (S80)
corresponding to the vacuum and the first excited state in the NS sector, respectively.
The crosscap Ishibashi states associated with the primary fields $\mathbbm{1}$ and $\varepsilon$ are given by [21]
$|\alpha\rangle\rangle_{\mathcal{C}}=\prod_{n=1}^{\infty}\left[\sum_{p_{n}=0}^{\infty}(-1)^{np_{n}}L_{-n}^{p_{n}}\bar{L}_{-n}^{p_{n}}\right]|\alpha\rangle\,,\quad\alpha=\mathbbm{1},\varepsilon\,.$ (S81)
Utilizing the fermionic representation of the Virasoro generators [Eq. (S79)] and the associated primary states [Eq. (S80)], these Ishibashi states can be expanded in the fermionic basis as
$|\mathbbm{1}\rangle\rangle_{\mathcal{C}}$ $=\sum_{M\;\mathrm{even}}\sum_{0<n_{1}<\cdots<n_{M}}(-1)^{\sum_{\alpha=1}^{M}(n_{\alpha}-1/2)}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\prod_{\alpha=1}^{M}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}\,$,
$|\varepsilon\rangle\rangle_{\mathcal{C}}$ $=-i\sum_{M\;\mathrm{odd}}\sum_{0<n_{1}<\cdots<n_{M}}(-1)^{-\frac{1}{2}+\sum_{\alpha=1}^{M}(n_{\alpha}-1/2)}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\prod_{\alpha=1}^{M}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}\,.$ (S82)
D.3 C. Continuum counterparts of the lattice crosscap states
The Ising CFT is the effective theory of the critical Ising chain [Eq. (S1) with $h=1$], where the lattice Majorana fermion operators $\chi_{j}$ and $\bar{\chi}_{j}$ in Eq. (S9) are identified with the Majorana fields $\chi(x)$ and $\bar{\chi}(x)$ in Eq. (S78), and the Bogouliubov modes near $\pm\pi$ are identified with the momentum modes of the Majorana fields:
@@
$|B_{\pm}\rangle=\exp\left[\pm\sum_{n=1}^{\infty}(-1)^{n}b_{-n+\frac{1}{2}}\bar{b}_{-n+\frac{1}{2}}\right]|0\rangle_{\mathrm{NS}}\,.$ (S86)
@@
To identify $|\mathcal{C}_{\pm}\rangle$ as the conformal crosscap states, we notice that the crosscap Ishibashi states $|\mathbbm{1}\rangle\rangle_{\mathcal{C}}$ and $|\varepsilon\rangle\rangle_{\mathcal{C}}$ [Eq. (S82)] can be expressed in terms of $|B_{\pm}\rangle$ [Eq. (S86)]:
$|\mathbbm{1}\rangle\rangle_{\mathcal{C}}$ $=\sum_{M\;\mathrm{even}}\sum_{0<n_{1}\cdots<n_{M}}(-1)^{\sum_{\alpha=1}^{M}(n_{\alpha}-1/2)}\,(-1)^{M/2}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}$
@@
$|\varepsilon\rangle\rangle_{\mathcal{C}}$ $=-i\sum_{M\;\mathrm{odd}}\sum_{0<n_{1}<\cdots<n_{M}}(-1)^{-1/2+\sum_{\alpha=1}^{M}(n_{\alpha}-1/2)}\,(-1)^{\frac{M-1}{2}}\prod_{\alpha=1}^{M}b_{-n_{\alpha}+1/2}\bar{b}_{-n_{\alpha}+1/2}|0\rangle_{\mathrm{NS}}$